OpenStudy (anonymous):

Complete the square: (x - 4)^2 = 49 (I'd like help, not answers) =)

OpenStudy (anonymous):

This is just a practice question, but I don't know how to do it and can't find anything in the lesson

OpenStudy (anonymous):

x = 12 x = 11 x = 4 x = -3 Those are the choices

OpenStudy (loser66):

49 = (what number)^2?

OpenStudy (anonymous):

7

OpenStudy (loser66):

good, so, (x-4)^2 = 7^2 see the outlet?

OpenStudy (anonymous):

Not really (sorry, I'm not good at math)

OpenStudy (loser66):

take off the square both sides, what are left?

OpenStudy (loser66):

(x -4) ^2 = 7 ^2 take off the square both sides

OpenStudy (anonymous):

(x-4) = 7 or (x-4)(x-4) = 49?

OpenStudy (loser66):

nope, take off x -4 = 7

OpenStudy (loser66):

now, +4 both sides

OpenStudy (anonymous):

x = 7 - 4?

OpenStudy (loser66):

x - 4 = 7 x -4 +4 = 7 +4 x = 11

OpenStudy (anonymous):

Okay. Sorry, wasn't paying attention to the sign. Makes sense. Thanks!